OpenStudy (anonymous):

could someone explain to me the laws of natural log e, example why is it that e^3In2 =8

OpenStudy (lgbasallote):

\[\huge e^{3\ln 2} \implies e^{\ln 2^3} \implies e^{\ln 8} \implies e^{\log_e 8} \implies 8\] does that help?

OpenStudy (anonymous):

yes thank you